### v2

# Analytical coarse grained potential parameterization by Reinforcement Learning for anisotropic cellulose

Abstract
Cellulose nanocrystals (CNC) are a type of cellulose with excellent mechanical performance and other merit attributes. According to previous reports, hydrogen bonds play a pivotal role in the anisotropic structure of the CNC. Understanding the structure and mechanical behavior of CNC on a mesoscopic scale is critical for the development and manufacture of cellulose materials. However, experimental observations and atomistic simulations are not appropriate on the mesoscopic scale. In this study, we introduce an analytical coarse-grained (CG) potential that is directly parameterized using Reinforcement Learning. Reinforcement Learning (RL) is a power source for industrial and academic applications in many fields. However, its potential has not yet been fully exploited in molecular dynamics. The RL and Boltzmann inversion methods were employed to develop a novel CG model of cellulose to represent its anisotropy and polymer stiffness. The resultant CG model is not limited to the target properties for training, and can present mechanical properties under other circumstances without additional training. This model confirms that RL can construct a CG potential that is both physically explainable and powerful. The training code is available at https://github.com/EiPiFun/rl-cll-cg.
Introduction
1.1 Cellulose nanocrystals
As the most abundant natural polymer resource on Earth, cellulose is widely available and eco-friendly, with many natural sources including wood, bamboo, algae, and bacteria[1, 2]. Renewability, low cost, biocompatibility, biodegradability, and good mechanical performance[3, 4] are the outstanding merits of cellulose, making it a good replacement for petroleum-based materials. As a nano-form of cellulose material, cellulose nanocrystals (CNC) can achieve superior performance, such as a high elastic modulus (approximately 200 GPa), low mass density (1.6 $\rm{g/cm^{3}}$), high specific surface area (several hundreds of $\rm{m_{2}/g}$), and a low coefficient of thermal expansion[4]. Owing to their excellent mechanical properties and interactions with other ions and polar molecules[5], CNC have been widely studied and employed in biomedicine[6], water treatment[7], energy storage[8], and humidity sensors[9]. Computational simulations, including ab initio simulations[10], molecular dynamics[11], and finite element method simulations[12], were performed to reveal the molecular details and kinetics of CNC. Experimental[13, 14, 15], computational[16, 17, 12], and analytical[18, 19] approaches have been exploited to understand the mechanical properties of CNC. Hydroxyls and HBonds play pivotal roles in these behaviors[20, 21, 22].
CNC are composed of parallel cellulose chains and their axial elastic modulus is much larger than their transverse modulus. The covalent bonds along the cellulose chains are much stronger than the van der Waals and HBonding (HBonding) interactions in the transverse section. The axial modulus and transverse modulus of CNC are approximately 100-200 GPa and 10-50 GPa, which is confirmed by experiments[4] and simulations[16, 17]. In the transverse section, the CNC are orthotropic and exhibit three characteristic load directions: along, perpendicular to, and slanted to the HBonding planes, as shown in Figure 1. Flat residues and HBonds are the key factors in the transverse anisotropy of cellulose. Wu et al. emphasized the anisotropy of CNC, and they simulated the stretch in three orthogonal directions (along the chain, parallel, and perpendicular to the HBonding planes) at three different strain rates and found that the mechanical properties (elastic modulus and Poisson’s ratio) are highly anisotropic and independent of the strain rate[17]. When a vertical load is applied, the slant CNC model undergoing frictional sliding is ductile, whereas the vertical and horizontal models are fragile. Frictional sliding is an important mechanical behavior of CNC and has been elucidated in previous studies. The friction of the nearby cellulose layers at different velocities, normal loads, and relative angles was also simulated. They found that the number of HBonds and the distances between layers were important factors in this case[23]. Zhang et al. systematically simulated and examined the impact of the loading direction, interfacial moisture, and misalignment on the CNC-CNC interface friction mechanical behavior. The density of HBonds is strongly correlated with interaction energy in their results[24]. The frictional sliding of CNC is critical to their mechanical performance when the loading direction is slanted toward the HBonding planes.
Refer to captionRefer to caption
Vertical
Refer to captionRefer to caption
Horizontal
Refer to captionRefer to caption
Slant
Figure 1: Orthotropic transverse section and characteristic directions. (a) Characteristic directions of orthotropic transverse sections dominated by HBonding planes. (b) Behaviors in characteristic directions. Flat residues and HBonds are critical for cellulose anisotropy. The vertical and horizontal models are dominated by HBonds and van der Waals interactions, resulting in fragile performances. The load of the slant model induced friction and rotation of cellulose residues.
To reflect the performance of CNC on a larger length scale, many coarse-grained (CG) models have been developed with different CG levels for different purposes such as mechanical properties and molecular interactions[25]. Phenomenon such as self-assembly or cellulose nano networks require supra CG models, which means that several cellulose residues become one bead[26, 27, 28] If the mechanical properties of the CNC are considered, one cellulose residue (one residue or two nearby residues in the same chain) is usually mapped to a bead[29, 30]. The directionality of the transverse sections was neglected. Some researchers have emphasized molecular details and interactions; thus, they will preserve more molecular details and map one cellulose residue to more than one bead instead of several residues into one bead[31, 32, 33, 34].
However, the CG models of cellulose mentioned previously could not reveal the HBond interactions and resultant directionality. Several studies have presented directionality using anisotropic particles[26, 35] or additional bonds that are not supported by real molecular structures to compensate for the directionality [36, 37, 38, 39]. Shishehbor et al. developed a CG CNC by matching the axial elastic modulus and transverse stress-strain curve to investigate the interface interaction and failure mechanism[38]. The CG mappings used by Shishehbor et al. neglected the flat structure of cellulose residue. Anisotropic particles and bonded (BD) directionality are unsuitable when a transverse failure mechanism with frictional sliding is considered in the CG model. Nonetheless, presenting the kinetics, energy, and mechanical properties of CG models is challenging[33].
1.2 Machine Learning parameterization and Reinforcement Learning
The data-driven Machine Learning potential may be the first consideration for a combination of Machine Learning and force fields. Machine learning models, particularly neural network models, can describe particle interactions without substantial physical or chemical intuition[40, 41]. However, significant problems remain for Machine Learning potentials, such as long-range interactions, transferability, and kinetics (particularly for CG models) are still open[40]. Neural network models also have problems with explainability[42] and computational cost. Machine learning methods can also be used to parameterize analytical force fields[43, 44]. Optimization methods, including Bayesian optimization[45], particle swarm optimization[46], genetic algorithm[47], and Gaussian process regression[48] have been used to parameterize different analytical CG potentials.
The use of Reinforcement Learning (RL) in CG simulations and potential parameterization is still in its early stages[49]. Wu et al. integrated graph attention neural networks and RL to learn CG mapping[50]. Chandra et al. used a continuous action Monte Carlo Tree Search[51] powered by RL to parameterize an analytical hybrid bond-order coarse-grained force field to simulate the aggregation of liquid-liquid mixtures[49]. RL imitates the learning of creatures and is promising for control and optimization[52, 53]. The model-free RL algorithm can deal with strongly nonlinear problems without prior expert knowledge and automatically perform trial-and-error learning procedures. More information on the RL scheme is presented in the methods section.
In summary, flat residues and HBond interactions are key features of CNC orthotropic transverse sections and their corresponding transverse mechanical behaviors, which have not been well described in previous CG models. With a tight budget for computational cost and a desire for better explainability, transferability, and kinetics, a carefully designed model with explicit analytical potential, instead of a data-driven potential model, was developed to reproduce the orthotropic mechanical behavior of the CNC. In this study, we introduce a novel CNC CG model parameterized by RL that considers flat residues, HBonds, polymer stiffness, and anisotropic mechanical performance. HBonds were implemented as nonbonded (NB) interactions with directionality in this model.
Methods
2.1 Mapping and topology
The mapping emphasizes the flat residues of the cellulose chains and the HBond interactions. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads, as shown in Figure 2,. Two features are considered: the distance between CL2 and CL3 in a residue and the angle of CL2-CL3 and the backbone. The first distance feature should be sufficiently long to bear the torque and present flat residues, whereas the second angle feature should be approximately 90. In the crystalline state, the angles of the nearby CL2:CL3-CL1(blue:orange-green) and CL3:CL2-CL1(orange:blue-green) (colons indicate that these two beads are not bonded together, whereas dash indicates bonded) are close to $180^{\circ}$, as illustrated in Figure 2.
Bead
Bond
Angle
Improper
Figure 2: Mapping and topology. (a) Mapping and (b) mapped overlays for the cellulose residue and (c) laminar HBonding plane. One cellulose residue was mapped to one backbone CL1 bead and two branched CL2 and CL3 beads. The basic idea behind this mapping is to preserve the flat structure and HBonds within the HBonding layer. Qualitatively, the distance between CL2 and CL3 from the same residue was sufficient and the branched beads were expected to be perpendicular to the backbones. Under this mapping, a layer of CNC was mapped as a network structure, where the HBond interactions were stressed. (d) BD topology. BD interactions are required to preserve molecular structure and performance under stretching, bending, and twisting loads. (e) NB topology. For HBond (HB) potentials, two symmetric A:H-D relations (CL2:CL3-CL1 and CL3:CL2-CL1) were defined to represent HBonds. HB potentials were determined using both the A:D distance and A:H-D angle.
Based on this mapping, we applied three types of bonds, six types of angles, and three types of improper dihedrals as the BD interactions (Figure 2). The BD interactions were designed to simulate and restrain the deformations under stretching, bending, and twisting loads. To ensure restraint in the aperiodic chain (which was used for the polymer stiffness simulations during training), the CL2-CL1-CL1 and CL3-CL1-CL1 angles were categorized. A type of directional NB interaction was introduced to describe HBonds, which requires the Acceptor, Hydrogen and Donor beads. Two A:H-D topologies were determined based on the importance of HBonding and symmetries: CL2:CL3-CL1 and CL3:CL2-CL1. This directional interaction is affected by the A:H-D angle and A-D distance[54].
2.2 Potentials
In this model, the potential energy contained both BD and NB interactions.
$V_{System}=V_{Bond}+V_{Angle}+V_{Improper}+V_{non_{b}onded}$ (1)
. The BD interactions were harmonic.
$V_{Bond}=\sum k_{B,i}(r-r_{0,i})^{2}$ (2)
$V_{Angle}=\sum k_{A,i}(\theta-\theta_{0,i})^{2}$ (3)
$V_{Improper}=\sum k_{I,i}(\chi-\chi_{0,i})^{2}$ (4)
. The NB interaction contained Lennard-Jones (LJ) and HBond (HB) potentials[55].
$V_{non_{b}onded}=\sum V_{LJ,i}+\sum V_{HB,i}$ (5)
in which $V_{LJ}$ and $V_{HB}$ were in the forms
$V_{LJ}=\cases{\epsilon_{LJ}[4{(\frac{\sigma_{LJ}}{r})}^{12}-4{(\frac{\sigma_{LJ}}{r})}^{6}]+V_{shift}&, $0<r\leq r_{in}$\cr\int_{r_{in}}^{r}[C_{3}(x-r_{in})^{3}+C_{2}(x-r_{in})^{2}+C_{1}(x-r_{in})+C_{0}]dx+V_{0}&, $r_{in}<r\leq r_{c}$\cr}$ (6)
$V_{HB}=\cases{\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $0<r\leq r_{in}\,\&\theta-\theta_{0}+\pi\geq\theta_{c}$\cr S(i)\cdot\epsilon_{HB}[5{(\frac{\sigma_{HB}}{r})}^{12}-6{(\frac{\sigma_{HB}}{r})}^{10}]{{\rm cos}^{n}(\theta-\theta_{0}+\pi)}&, $r_{in}<r\leq r_{out}\&\theta-\theta_{0}+\pi\geq\theta_{c}$\cr}$ (7)
where:
$S(i)=\frac{{(r_{out}^{2}-r^{2})}^{2}[r_{out}^{2}+2r^{2}-3r_{in}^{2}]}{{(r_{out}^{2}-r_{in}^{2})}^{3}}$ (8)
The LJ potential is a 12-6 formula of Lennard-Jones interactions, and the HB potential is a 12-10 formula with scaling by the power of the cosine of the A:H-D angle. Technically, these are LAMMPS[54] pair styles. The LJ potential is the lj/smooth with a shift, whereas the HB potential is a modified version of the hbond/dreiding/lj. lj/smooth is designed to smooth force to zero at the cutoff radius using a polylinear expression; this modified version of hbond/dreiding/lj ensures that potential energy reach minima at $\theta_{0}$ instead of $\pi$. The energy and force curves of the HB and LJ potentials are presented in Figure SP9.
The equilibrium A:H-D angles in the CG mapping were designed to be close to $\pi$. Even though we generalize it by adding an equilibrium angle of A:H-D, an equilibrium angle close to $\pi$ is suggested to stabilize the conical energy surface.
2.3 Boltzmann inversion for harmonic potential
Boltzmann inversion is structure-based and is widely utilized to determine bonded potentials. Boltzmann inversion can also be integrated with Machine Learning methods to construct potentials[46, 56] or used as a baseline[57, 58].
For a harmonic potential following the Boltzmann distribution, where the bond length (or angle, or dihedral) is a deviated $x$ around the equilibrium length $x_{0}$:
$\rho(x+x_{0})\propto e^{-\frac{V}{k_{B}T}}=e^{-\frac{V_{x}-V_{0}}{k_{B}T}}=e^{-\frac{kx^{2}}{2k_{B}T}}$ (9)
. Mark $\rho(x_{0})=\rho_{0}$:
${\rm ln}\rho(x+x_{0})=-\frac{kx^{2}}{2k_{B}T}+{\rm ln}\rho_{0}$
$k=-2k_{B}T\cdot\frac{d{[{\rm ln}\rho(x+x_{0})]}}{d{[x^{2}]}}$ (10)
. We then mark $\frac{d{[{\rm ln}\rho(x+x_{0})]}}{d{[x^{2}]}}=\kappa$:
$k=-2k_{B}T\cdot\kappa$ (11)
However, the $k$ derived must be reparameterized to represent the mechanical performance[59, 60, 49] of the coarse-grained models with fewer degrees of freedom. The effective $k$ is $k_{0}$ times of $\kappa$:
$k_{i}=k_{0}\kappa_{i}$ (12)
. Based on $\kappa$ values, three $k$ values were used for the bond, angle, and improper force constants. Considering the axial elastic modulus and polymer stiffness, $k_{B0}$, $k_{A0}$, and $k_{I0}$ were determined via RL optimization.
2.4 Reinforcement Learning and degenerate RL
In RL, the entire problem is called the environment and the optimization program is called the agent. The environment provides at least three types of information: observation, action space, and reward. In an interaction episode with the environment, the agent selects an action $a_{t}$ in the action space based on the obtained partial observation $o_{t}$ of the state of the environment. The action of the agent impacts the environment evolution and is evaluated using the reward $r_{t}$ function. By continuously interacting with the environment, the agent gradually learns to perform actions corresponding to a given observation to maximize the cumulative reward. The cumulative reward is defined as
$R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$ (13)
with $\gamma\in(0,1)$. The discount factor $\gamma\in(0,1)$ emphasizes the latest episode and ensures numerical convergence.
Figure 3: RL diagrams. (a) Standard RL diagram. In the standard RL diagram, the agent interacts with the environment continuously and performs actions corresponding to observations. Its optimization target is to maximize the cumulative reward $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. (b) One-shot RL diagram. However, a one-shot diagram was applied and only one episode was processed. In this study, RL was employed as a nonlinear optimizer, guided by its reward function.
In this study, a degenerate version of RL was leveraged in a one-shot diagram. There is only one step in a learning episode, and only one reward is sampled, because the performance of a set of coefficients is independent of that of the others. Therefore, the reward can be simplified as
$R(\tau)=r_{0}$ (14)
. Consequently, the RL algorithm was exploited as a nonlinear optimizer, guided under indirect supervision by a reward function. The use of degenerate RL was inspired by the work of Viquerat et al.[61], which is rare, particularly for potential parameterizations.
The RL agent used in this study was driven by the Soft Actor Critic (SAC) algorithm[62]. SAC is the state-of-the-art algorithm recommended for continuous action space by the doc of OpenAI Stable-Baselines3[63]. The key feature is the reconciliation between the expected reward and entropy, which implies that stochasticity is included in the SAC policy to a certain extent. SAC is a gradient-free and model-free method that does not rely on a specific problem environment, and can address strongly nonlinear problems after sufficient training. Detailed information on the SAC algorithm can be obtained in the doc of OpenAI Stable-Baselines3[63].
During training, the action consists of 17 normalized coefficients ranging from [-1,1] because most RL algorithms rely on a standard normal distribution[63]. actions will be transformed linearly into physical space corresponding to 3 BD $k$, 1 HB $\epsilon$, 1 HB $\sigma$, 6 LJ $\epsilon$ and 6 LJ $\sigma$. The reward function is a critical factor that guides the learning of the agent, and is defined by a series of matching degrees. The matching degree $m$ for a specific performance is defined as
$m=\frac{1}{1+|\frac{y_{1}-y_{0}}{y_{0}}|}$ (15)
with a threshold $m_{t}$
$M=\cases{m&, $m\geq m_{t}$\cr 0&, $m<m_{t}$}$ (16)
. where $y_{0}$ and $y_{1}$ are the reference and sampled values, respectively. The reward function is the sum of a series of matching degrees, except for the transverse strength $M_{s}$ and toughness $M_{t}$,
$reward=\sum M_{i}+w\cdot{\rm min}(M_{s},M_{t})$ (17)
because transverse strength and toughness were the hardest to match during the training procedures. In RL, the reward functions must be accurate and sufficiently discriminant[64]. The weights of $M_{s}$$M_{t}$ and the thresholds were designed to enhance discrimination.
2.5 Simulation setup
For both all atom (AA) and coarse-grained (CG) simulations, three models in different arrangement directions under the NPT ensemble at 0.1 MPa with a stretch load in the vertical direction were simulated. All reference AA simulations were performed using GROMACS[65, 66] and CHARMM36[67, 68] force field (force field files were generated using CHARMM-GUI tools[69, 70]). All the CG simulations were performed using LAMMPS[54] with a self-defined force field. Referring to the performance-size curves obtained by Sinko et al.[71], we used a CNC model comprising 144 cellulose chains. Fragile vertical and horizontal models with 72 cellulose chains were used to reduce the computational cost of the simulations during training. Vertical stretches were implemented using cell deformation loads at a constant speed 10.0 nm/ns. 100 replica AA simulations with different stretch speed ranging from 0.1 nm/ns to 40.0 nm/ns are performed. The AA simulation results confirmed that the sampled strength and toughness were stable within this speed range (with deviations of less than 15%), as shown in Figure SP9. Figure LABEL:fig:aa_transverse_stretch_performance_collection shows the stress-strain curves of the 100 replica simulations.
2.6 Training procedures and statistics-guided reduction
Before training, the equilibrium geometry parameters (bond length $l^{e}$/,angle size $\theta^{e}$/, and dihedral size $\chi^{e}$) were derived from the mapped AA trajectories. The estimated force constants for the bonded harmonic interactions were determined using the same trajectories and Boltzmann inversion method. Three rescale factors were required for the bonds, angles, and improper dihedrals in this model. Although rescaled, the relative ratios of the force constants were maintained (e.g., the force constants for improper dihedrals were rescaled with the same factor). As previously mentioned, there are 17 coefficients, where three are BD coefficients and 14 are NB coefficients. To simplify the problem and reduce computational cost, three training stages were designed (Figure 4). The first stage was BD training for three BD $k$ only, ignoring all NB interactions and targeting the BD properties (axial elastic modulus and polymer stiffness). We cannot determine the influence of NB interactions in the first stage, and we assume that they are not critical for the BD properties. The second stage was NB training for 14 NB coefficients with the BD coefficients obtained from the first stage, targeting the BD and NB properties (transverse strength and toughness in the three characteristic directions). The third stage was preserved only if the performance in the second stage was unsatisfactory; all properties were targeted to parameterize 17 coefficients at every turn.
The training results indicated that NB interactions did not play a pivotal role in the target BD properties. Therefore, our training ended in the second stage.
Figure 4: Training procedures. The BD properties (axial elastic modulus and polymer stiffness) and NB properties (transverse strength and toughness in characteristic directions) were the target attributes. Before training, the equilibrium bonded geometry parameters and force constants of the BD harmonic potentials were estimated from the mapped AA trajectories using the Boltzmann inversion method. In the first stage, the force constants were scaled by the RL agent in the BD training to reproduce the BD properties. In BD training, all NB interactions were ignored because we assumed that the NB interactions were not significant in this case. Given both BD and NB properties, NB training using the previously obtained BD coefficients was performed to match both BD and NB properties. However, only one batch of simulations was performed for the target properties during training. Thus, the transverse stretch performance data were manually verified through replica simulations to confirm the statistical agreement. Multiple sets of coefficients were collected to reduce the number of independent coefficients for the computational cost and physical explainability. The results confirm that our assumption of the BD properties is reasonable. Therefore, the coupled training was unnecessary.
However, there was also important artificial intervention during the training process. The transverse stretch performance was sampled only once to control the cost, which was insufficient to ensure accuracy of the distribution levels between the AA and CG models. Therefore, manual tests for the optimized coefficients were required to verify the performance distribution, and the optimized coefficients were analyzed, inspired by Befort et al.[44]. Only coefficients with significant tendencies rather than those without clear expectations were retained as independent coefficients. These statistics-guided reduction procedures are described in the results section.
Results
3.1 Coefficients
The distance coefficients ($\sigma$) of the NB interactions were estimated using the equilibrium distance of the mapped AA trajectories. The equilibrium distance $d^{e}$ is defined as the average of the minimum distances for the NB bead pairs.
$d^{e}=<d^{e}_{t}>=<d^{e}_{tm}>=<<{\rm min}(d^{e}_{tmn})>>$ (18)
[…]
Figure 5: Training convergence and statistics-guided reduction. (a) Reward and its moving average in the first and (b) last training loops. Each training loop contained 1024 training steps. The subsequent training loops with higher and more stable rewards were based on former training loops. (c) Average and maximum rewards for different training loops. The average and maximum rewards confirmed the convergence. After training, the best coefficients for the different training loops were collected to perform statistics-guided reduction. The first reduction was inferred from the geometric characteristics of the CG models, and a constraint of $\sigma_{p11}=\sigma_{p22}=\sigma_{p33}$ was applied (pmn presents a pair of CLm and CLn), which is also helpful for fragile behaviors in the vertical and horizontal directions. (d) The second reduction was in the distance coefficients using statistics. $\sigma$ except for $\sigma_{p23}$ oscillated around its equilibrium distance $d^{e}$ whereas an obvious increase of $\sigma_{p23}$ was observed. Then only $\sigma_{p11}$ and $\sigma_{p23}$ are kept as independent distance coefficients: $\sigma_{p12}=\frac{d^{e}_{p12}}{d^{e}_{p11}}\sigma_{p11},\sigma_{p13}=\frac{d^{e}_{p13}}{d^{e}_{p11}}\sigma_{p11}$. (e) The third reduction in energy coefficients was statistically determined. After the first and second reduction, we observe that $\epsilon_{p23}$ is usually the smallest energy coefficients, which may correspond to the $\sigma{p23}$ enlargement. Then only $\epsilon_{p23}$ and $\epsilon_{p23}$ are kept as independent energy coefficient: $\epsilon_{p11}=\epsilon_{p22}=\epsilon_{p33}=\epsilon_{p12}=\epsilon_{p13}$.
The coefficients were trained using several training loops, each containing 1024 training steps. When a training loop was completed, the information of the agent was saved and loaded for subsequent training. Referring to the first training loop, the last training loop achieved a higher reward much faster and more stably (Figure 5a). The average and maximum rewards of the different training loops were compared (Figure 5b) to confirm the convergence of training.
When the training was confirmed to be convergent, the coefficients with the highest reward in each training loop were extracted to determine important coefficients. The first reduction in the NB coefficients was inspired by the laminar geometry.
$\sigma_{p11}=\sigma_{p22}=\sigma_{p33}$ (19)
where pmnpmn is the pair of the CLm and CLn. The structures of the CG models with lower degrees of freedom were more orderly. Thus, the equilibrium distances of the same types should be close in order to represent the laminar structure. In other words, the equilibrium distance of CL1-CL1, CL2-CL2 and CL3-CL3 ought to be nearly the same, whereas in the mapped AA trajectories, they are not (Figure SP9). The equilibriumequilibrium distancesdistances of the ideal crystal structure were selected as the estimated values of the distance coefficients ($\sigma$) for the pairs. These equilibriumequilibrium distancesdistances do not correspond to the well-relaxed structures or characteristic distances of the radial distribution function. The first physics-informed restraint was also beneficial for reproducing fragile fractures in vertical and horizontal models.
The subsequent reductions in the NB coefficients were statistically guided.
Second reduction: Referring to $d^{e}$, only $\sigma_{p23}$ had an unexpected intention to increase, maintaining $\sigma_{p11}$ and $\sigma_{p23}$ only.
$\sigma_{p12}=\frac{d^{e}_{p12}}{d^{e}_{p11}}\sigma_{p11},\sigma_{p23}=\frac{d^{e}_{p23}}{d^{e}_{p11}}\sigma_{p11},\sigma_{p13}=\frac{d^{e}_{p13}}{d^{e}_{p11}}\sigma_{p11}$ (20)
Third reduction: $\epsilon_{p23}$ were usually the smallest energy coefficients in the system, trying to keep only $\epsilon_{p11}$ and $\epsilon_{p23}$ independent
$\epsilon_{p11}=\epsilon_{p22}=\epsilon_{p33}=\epsilon_{p12}=\epsilon_{p13}$ (21)
After these dimensionality reduction, only 9 independent coefficients are preserved: 3 BD $k$, 1 HB $\epsilon$, 1 HB $\sigma$, 2 LJ $\epsilon$ and 2 LJ $\sigma$.
The determined BD coefficients are listed in Table 1 and the NB coefficients are listed in Table 2. All coefficients are given with four significant digits.
Table 1: Bonded coefficients
[…]
bond 1 64.06 5.270 5.270 5.258
2 82.42 2.493 2.493 2.473
3 112.3 2.066 2.066 2.061
kk(kcal/mol) $\theta_{0}(^{\circ})$ $\theta^{e}_{\rm AA}(^{\circ}))$ $\theta^{e}_{\rm CG}(^{\circ}))$
angle 1 30.97 163.5 163.5 162.3
2 38.79 74.2 74.2 74.7
3 30.52 93.2 93.2 94.0
4 27.28 90.0 90.0 89.2
5 25.97 102.0 102.0 101.4
6 22.42 166.3 166.3 165.1
kk(kcal/mol) $\chi_{0}(^{\circ})$ $\chi^{e}_{\rm AA}(^{\circ}))$ $\chi^{e}_{\rm CG}(^{\circ}))$
improper 1 0.1702 178.0 178.0 175.6
2 0.1555 2.3 2.3 2.1
3 0.1698 2.9 2.9 2.1
Table 2: Nonbonded coefficients
type pair $\epsilon$(kcal/mol/Å) $\sigma$(Å) $d^{e}_{\rm AA}$(Å) $d^{e}_{\rm CG}$(Å)
HB 2:3-1 1.404 6.017
3:2-1 1.404 6.017
LJ 1-1 0.9031 5.101 5.398 5.750
1-2 0.9031 4.201 4.769 4.473
1-3 0.9031 4.101 4.712 4.609
2-2 0.9031 5.101 4.636 5.260
2-3 0.5031 4.083 3.983 4.018
3-3 0.9031 5.101 4.528 5.271
3.2 Bonded properties
For the BD interactions and validation of the BD coefficients, we considered the axial elastic modulus and polymer stiffness of a single cellulose chain, as listed in Table 3 and shown in Figure 6. Axial stretching was performed in an xyz-periodic cell with a deformation load along the chains, which is a basic mechanical property. The simulation of polymer stiffness is a single chain centered in a cell. For axial stretch, Poisson effect was presented in CG model and elastic modulus of AA and CG are 133.5 GPa and 130.3 GPa respectively, where reference modulus is 120-140 GPa[13, 14] from experiments and 110-200 GPa[16, 38, 17] from simulations. For polymer stiffness, persistence length of AA and CG range from 9.7-9.8 nm and 8.9-10.2 nm with reference data of 6-25 nm from experiments[72] and simulation[73]. Notably, the polymer stiffness data were not stable for the AA simulations because of the intertwining in some cases. Therefore, persistence lengths and end-to-end distances of different lengths (10, 20, and 30 residues) are listed. When considering polymer stiffness, the joint O atoms between two cellulose residues in the AA models and CL1 beads on the backbone in the CG models were node particles to be considered. The reference data for training were based on AA information only, and the mapped AA trajectories were compared with the CG trajectories after training. In this study, all effective $k$ values were smaller than their estimations based on distributions (Table SP2), particularly for angles and dihedrals, to reproduce the polymer stiffness performance in a system with fewer degrees of freedom (Table SP2). A comparison of the bonded geometry parameter distributions is shown in Figure SP9.
Table 3: Bonded properties
type AA CG Error
Axial modulus (GPa) 133.5 130.5 2.25%
Persistence
length (nm)
10 9.755 8.990 7.84%
20 9.815 10.176 3.68%
30 9.716 10.106 4.01%
End-to-end
distance (nm)
10 4.523 4.529 0.01%
20 9.355 9.535 3.52%
30 13.536 13.536 0.08%
Table 4: Nonbonded properties
type AA CG Error
Strength
[…]
Toughness
[…]
Direction
angle
[…]
Figure 6: Bonded properties. (a) Axial elastic modulus. The elastic modulus of AA and CG are 133.5 GPa and 130.5 GPa separately, corresponding with each other and reference[38]. (b) Polymer stiffness data containing persistence length and (c) end-to-end distance. The persistence lengths of the AA and CG models with three different lengths were approximately 10 nm with errors of less than 15% and in agreement with a previous study[72]. The end-to-end distances of the AA and CG models were also similar, with errors of less than 5%.
3.3 Nonbonded properties
The NB properties include the transverse strength and toughness, as shown in Figure 7. Table 4 lists the NB properties numerically. The reference strengths of the vertical, horizontal and slant models are 0.9 GPa[17], 0.5 GPa[17] and 0.7 GPa[38], which were well reproduced by both AA and CG simulations. To reduce the effect of randomness, 100 replica simulations of transverse stretch were performed. The strength and toughness distributions of the CG models were similar to those of the AA reference data. Only the distributions of the horizontal CG models were narrower than those of their AA counterparts were. The average strength and toughness values were similar to those obtained from the AA data. The frictional sliding of the slant models was quantitatively validated by using the direction angle (Figure SP9). These data validate the CG model for reproducing the transverse anisotropy and frictional sliding. Although all $\sigma$ values were larger than expected, the equilibrium distances in the CG models were only slightly larger than those in the AA models (less than %15, Table 2). A comparison of the nonbonded equilibrium distance distributions is shown in Figure SP9. The stretch curves (Figure SP9), radical distribution function, radius of gyration, and root-mean-square deviation were reproduced, as shown in Figure LABEL:fig:structural_properties.
Refer to captionRefer to caption
Vertical
Refer to captionRefer to caption
Horizontal
Refer to captionRefer to caption
Slant
Vertical
Horizontal
Slant
Vertical
Horizontal
Slant
Figure 7: Nonbonded properties. (a) Behaviors of CG models in the characteristic directions. Important frictional sliding in the slant direction and fragile fractures in the vertical and horizontal directions are reproduced. (b) Strength and (c) toughness distribution of transverse stretches. A total of 100 replica simulations were performed with different random seeds for both AA and CG models in three directions. For the vertical and slant models, the strength and toughness histograms were reproduced consistently, because their mean values and standard errors were similar. Despite the accuracy of the mean values, the horizontal CG models provided narrower distributions, which was within expectations.
3.4 Additional inspection
Additional inspections of the mechanical properties of the CG model were performed to verify its transferability and generalization to other cases.
3.4.1 Draw-out and Tear-apart
Opening, in-plane shear, and out-of-plane shear are the three types of fracture. The transverse stretch used in training corresponded to the type of opening fracture. For in-plane and out-of-plane shear, the draw-out and tear-apart were tested using steered molecular dynamics(Figure 8). A draw-out test was performed in a periodic cell under the NPT ensemble, and the middle cellulose chain was pulled out of the CNC. A tear-apart test towed the tips of the middle part of the cellulose chains perpendicular to their backbones under the NVT ensemble. The structure and force-displacement curves are shown in Figure 8. Determined by the periodic structure and HBonds of cellulose, draw-out was a periodic process with force fading. The tear-apart test was a type of force-increasing process, and the displacements and maximum forces of the AA and CG simulations are similar. These results indicate that the CG model can accurately illustrate the three types of CNC fracture.
Refer to captionRefer to captionRefer to caption
AA
Refer to captionRefer to captionRefer to caption
CG
AA
CG
Refer to captionRefer to caption
AA front view
Refer to captionRefer to caption
AA top view
Refer to captionRefer to caption
CG front view
Refer to captionRefer to caption
CG top view
Figure 8: Draw-out and Tear-apart. (a) Draw-out and (b) curves. (c) Tear-apart and (d) curves. The transverse stretch resembled an opening fracture, whereas the in-plane and out-of-plane shear were examined using draw-out and tear-apart simulations. The draw-out and tear-apart were performed separately under the NPT and NVT ensembles by steered molecular dynamics with axial towing of the middle cellulose chain or perpendicular towing of the middle chains. The force-displacement curve of the draw-out presented periodic fading force curves that were ruled by their structure, whereas the tear-apart force was incremental. The maximum force and frequency of the draw-out and tear-apart matched those of the AA references. Therefore, the CG model can quantitatively represent the in-plane and out-of-plane shear.
3.4.2 Adhesion and Bending
Although our CG model was designed for pure periodic systems, adhesion and bending loads were applied to verify its applicability to aperiodic systems (Figure 9). An adhesion test was performed to determine the adhesion energy of nearby aperiodic CNC in the axial direction. A vertical load was applied at the center point of the cellulose bundle for the bending test. The adhesion energy of the CG model was evidently larger than that of AA, and the slightly lower force of the CG model in bending may result from the BD force constant rescaling of the angles and dihedrals. The results confirmed that the CG model could also be used in aperiodic CNC systems with slightly lower accuracy.
Refer to captionRefer to caption
AA
Refer to captionRefer to caption
CG
Refer to captionRefer to caption
AA
Refer to captionRefer to caption
CG
Figure 9: Adhesion and Bending in aperiodic systems. (a) Adhesion. Two CNC were placed tightly close along the chain direction and were clung by the interface adhesion energy. They were then torn away from one another. The adhesion energy of the CG model was larger. (b) Bending. A vertical load was applied at the center of the CNC bundle to induce bending. The force curves of adhesion and bending confirm the capabilities of the CG model in aperiodic systems. Unlike the axial modulus dominated by bonds, the bending is also affected by the angles and dihedrals. The slightly lower bending modulus was assumed to be caused by the rescaling of the angle and dihedral force constants.
3.4.3 Transverse arrangement
Unlike previous BD directionality CG models[36, 38, 37, 39], the CG model was designed to present directionality via branched beads and HBond-like interactions. This makes it possible to consider the transverse arrangements of CNC with cross and symmetric patterns. The results of the arrangement patterns confirmed that the fractures and performance differed significantly from those of the crystals, as shown in Figure 10. The symmetric pattern was inspired by the twin structure, and the results indicated that this pattern exhibited a better toughness than the crystal. The cross pattern was symmetric in both directions. In these two patterns, frictional sliding was impeded and transformed into local rotations, thereby achieving better toughness and ductility in a symmetric pattern. These results emphasize the importance of the transverse arrangement, which cannot be described by former coarse-grained models.
Crystal
Symmetric
Cross
Refer to captionRefer to caption
Crystal AA and CG
Refer to captionRefer to caption
Symmetric AA and CG
Refer to captionRefer to caption
Cross AA and CG
Crystal
Symmetric
Cross
Figure 10: Transverse arrangement. (a) Transverse arrangement patterns. The crystal pattern was the same as that used for the potential parameterization. The symmetric pattern was inspired by the twin structure. The cross pattern was an X-shaped arrangement. (b) AA (left) and CG (right) structures under same vertical strain. In the symmetric and cross patterns, the global stretch was composed of the deformation of units, whose frictional sliding was hampered. In this case, local rotations became the dominant behavior, leading to better toughness and ductility. (c) Stress-strain curves of the three patterns. The transverse arrangements in the symmetric and cross patterns cannot be expressed by previous CG models. These arrangements should be emphasized in future studies.
3.4.4 Brick-and-mortar
[…]
AA
Refer to caption Refer to caption
CG
Figure 11: Brick-and-mortar. (a) Brick-and-mortar models and fractures. (b) Stress-strain curves when the elementary cellulose chain length was 40 residues. (c) Strength and (d) toughness curves of different elementary chain lengths. Highly aligned CNC can assemble into brick-and-mortar structures [74, 75] in the axial direction, and this staggered structure can improve strength and toughness. The brick-and-mortar stretch was also a CNC-level in-plane shear, and was heavily influenced by the length of the elementary cellulose chains. The CG model accurately describes the structural behavior of the brick-and-mortar. The brick-and-mortar fracture was composed of major local slips and decentralized minor slips, and the toughness is positively correlated with the overlapping length.
3.4.5 Computational efficiency
The computational efficiencies of the AA and CG models were compared using equilibrium simulations. A sequence of square periodic CNC (the word square indicates that the row and column numbers of the transverse section of the CNC were the same) equilibrium simulations of different sizes were used, in which the atom number of the AA models ranged from seven kilos to one million. The benchmark simulations were performed using central processing units (CPU) and graphical processing units (GPU). The HB interaction does not support hardware acceleration, and the AA simulations using only the same amount of CPU resources are shown in Figure 12. For the AA and CG simulations using only the CPU, six cores were allocated exclusively to each task. Six cores and an additional exclusive GPU were allocated for the accelerated AA simulations using a GPU. The results confirmed that the CG model was 20 times faster than its AA counterpart when using the same resources. Even when referring to the AA models with more acceleration, the efficiencies CG models were triple. When a GPU-accelerated HB interaction is implemented and tested in the future, its potential can be fully exploited.
Figure 12: Computation efficiency data (a) and (b) relative efficiency data with AA simulations as a baseline. Currently, the HB interactions used does not support hardware acceleration. Referring to the AA simulations with same amount of CPU resources, the CG model was 20 times faster in CNC equilibrium simulations of different transverse sizes. The efficiency of the CG model was still triple when additional exclusive GPU resources were utilized for AA. However, its computational efficiency has not yet been fully proven.
Discussions
Cellulose materials are usually composed of long and thin cellulose nanofibrils, and previous CG models of cellulose have emphasized their axial properties[29, 30]. However, CNC are anisotropic and much shorter, and their transverse properties should not be disparaged, which is important for future development of cellulose materials. Some exhibit transverse anisotropy using anisotropic particles[26, 35] or additional artificial bonds[36, 37, 38, 39], which are still incapable of representing the transverse anisotropy.
The challenges and main characteristics of the CG model include the flat residues and HBond (HBond) interactions. The CG model could reveal the frictional sliding and transverse arrangement, which are two critical properties for CNCs transverse anisotropy, and the former CG models could not reflect accurately. Except for the HBond interactions, only the harmonic and Lennard-Jones potentials were included in the BD and NB interactions. This model also confirms the potential parameterization capability of RL; only a discriminant reward function was required. The reward function for RL includes the matching degree of the BD properties (axial modulus and polymer stiffness) and the NB properties (transverse strength and toughness). The resultant model can reproduce the xyz three-dimensional anisotropy and transverse frictional sliding. RL may be more widely applied to potential parameterizations in the future.
However, several problems are encountered in the CG model. The first is he rescale of the angle and dihedral interaction force constants. The CG model, which has significantly fewer degrees of freedom, must compensate for the polymer stiffness properties by decreasing the force constants (Table SP2). The former CG model addressing persistence length did not require a precise match (40 nm)[34]. Some studies also estimated the persistence length of a bundle of cellulose chains instead of a single free chain[76]. These smaller force constants may cause problems for other properties involving the angles and dihedrals. For example, a lower stress was observed during the bending test.
The nonbonded equilibrium distances are also problematic. The equilibrium distances of the CG model were larger than those of their AA counterparts (Figure SP9), and the enlargement of the CL2-CL3 NB distance coefficients was determined using RL (Figure 5). This means that the distance coefficients and equilibrium distances do not correspond to the well-relaxed AA equilibrium distances or radial distribution function. In this study, the equilibrium distance of a pair was defined as the average minimum distance. Nearly all minimum distances of pairs are from beads of nearby HBonding layers; only the minimum distances of CL2-CL3 are from beads in the same HBonding layer. Therefore, we infer that the equilibrium distances from nearby HBonding plane pairs are appropriate estimations of $\sigma$ values for the LJ potentials because the dominant NB interactions are HBonds in the HBonding layers. The overall enlargement of the NB distance coefficients is assumed to be required by the increased $\sigma_{p23}$ to retain the crystal structure. Moreover, the enlargement of $\sigma_{p23}$ may be an assimilation of $\sigma$ because the equilibrium distance between CL2 and CL3 is the smallest. The Phenomenon of frictional sliding may be presented more easily if all $\sigma$ values are close to each other.
Other minor problems also exist. Our model may not perform well in aperiodic cellulose systems (as in the adhesion test). The pressure curve of the CG model under the NPT ensemble exhibited a larger oscillation amplitude. Otherwise, some molecular details were ignored in the CG model. For instance, the rotational degrees of freedom of the hydrogen atoms on the branch hydroxyls were neglected in harmonic BD potentials (Figure SP9). Moreover, one technical problem temporarily restrained the application potential of the CG model: the HBond-like pair style does not support hardware acceleration. Therefore, the computational efficiency of the CG model has not yet been fully realized.
As an early bird that introduced HBonds, the CG model of CNC could promote further research. The development of cellulose CG models has mainly focused on the native cellulose. However, solvent environments, particularly humidity, are key factors for cellulose materials[34, 77, 78, 79] and are not well represented in cellulose CG models[30]. Chemical surface modifications and interactions with other molecules such as hemicellulose or lignin are also critical for enhancing the performance of cellulose materials. These problems are tightly connected to the hydroxyls of cellulose residues. Regarding the computational resources and expertise, these interactions were not considered in this study. With the help of HBond-like interactions, a later CG model could attempt to present more properties of the cellulose. We hope that this model will facilitate future development of CG models.
As an application of RL for potential parameterization, our methods and implementations are not heavily dependent on cellulose, and may help in later CG model development. In the CG model, many priori knowledge including pre-defined mapping and interaction potential expressions are used to drop the difficulty and enhance physical explainability. However, the resultant effect of priori knowledge is not sure. Fully automatic optimization may achieve models with better performances. Transfer Learning and Imitation Learning methods may also be applied in the future to further accelerate RL in the potential parameterization applications.
Conclusions
In this study, we introduce a novel coarse-grained cellulose model with explicit branch beads and HBond-like interactions that can reproduce the transverse frictional sliding and anisotropic mechanical properties of cellulose nanocrystals. The strength and toughness in the three characteristic directions of the orthotropic transverse section were quantitatively reproduced without anisotropic particles or additional bonds. This study is an early bird that applied degenerate Reinforcement Learning as a nonlinear optimizer to perform direct analytical potential parameterization, and the physical meanings were enhanced by integrating the Boltzmann inversion method. This model is eclectic of physical knowledge, traditional modeling, and Machine Learning methods. It assists in the understanding of the importance of hydrogen bonds and the transverse anisotropy. With an appropriate reward function based on the target properties and statistics-guided reduction, the Reinforcement Learning agent can determine the appropriate coefficients without consuming considerable computational resources. The target anisotropic properties considered included the bonded properties (axial elastic modulus and polymer stiffness) and nonbonded properties (transverse stretch strength and toughness), all of which were accurately matched with errors below %15. Several additional inspections were performed after the training, including in-plane shear, out-of-plane shear, adhesion energy, bending, transverse arrangement, and brick-and-mortar structures. The transferability and generalization of this coarse-grained model were proven through additional inspections. These results emphasize that transverse anisotropy is important for transverse arrangements, which has not been quantitatively presented in previous coarse-grained models. Two transverse arrangement patterns with improved toughness and ductility were introduced. We found a positive correlation between the overlapping length and toughness of the brick-and-mortar structures. Data on transverse arrangements and brick-and-mortar structures may aid advanced material design in the future. This model may also aid in addressing interactions between cellulose and other polar molecules on a mesoscopic scale. This study emphasizes that Machine Learning can also be applied to construct effective analytical potentials that are better in terms of explainability, transferability, and kinetics. Although our results demonstrate the efficiency of Reinforcement Learning, many refinements remain to be explored, especially in the field of molecular research. For example, the necessity of inspiration from human knowledge and traditional methods such as Boltzmann inversion is still unclear. Finally, our basic parameterization approach is agnostic to the application case, implying that our workflow may also be applicable to other molecular systems. Transfer Learning and Imitation Learning methods may also be integrated with Reinforcement Learning to solve similar problems in the future.
Code availability
The modified hbond/dreiding/lj/angleoffset was merged in the official LAMMPS at
https://github.com/lammps/lammps/tree/develop/src/EXTRA-MOLECULE.
The code for potential parameterization training is available at
https://github.com/EiPiFun/rl-cll-cg.
Training data for the trained coefficients, pre-processing codes, and post-processing codes are available upon requests.
Conflict of Interest declaration
The author declares no competing financial interest.
Acknowledgments
Xu Don thanks the Department of Engineering Mechanics of Zhejiang University for the funding and computational resources.
Supplementary Information
The following subsections and data are included in the Supplementary Information.
Bonded properties without rescale
Force and energy curve of HB and LJ potentials
All atom transverse stretch performance at different stretch speed
Transverse stretch curves and distributions
Bonded geometry parameter distributions
Nonbonded equilibrium distance distributions
Other structural properties
Secondary summits
Table SP1:
Table SP2:
Figure SP1:
Figure SP2:
Figure SP3:
Figure SP4:
Figure SP5:
Figure SP6:
Figure SP7:
Figure SP8:
Figure SP9: